OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 4 9x – 3y = 14 A. There is no solution. B. (12, 14) C. (14, 12) D. There are an infinite number of solutions.

OpenStudy (anonymous):



OpenStudy (anonymous):

 take this and graph the two equations

OpenStudy (texaschic101):

y = 3x - 4 9x - 3y = 14 sub in 3x - 4 in for y in the 2nd equation 9x - 3(3x - 4) = 14 9x - 9x + 12 = 14 -- subtract 12 from each side 9x - 9x = 14 - 12 0 = 2 (incorrect) THERE IS NO SOLUTION